### v2

# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

Abstract
Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery. Mathematical competitions, in particular, presents two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation. We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving. We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations. On ConstructiveBench dataset, ECP improves the accuracy of the answer construction from Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with gpt-4.1-mini model. Moreover, combining with ECP’s constructed answers, state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines. Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
1University of Toronto  2Vector Institute  3Georgia Institute of Technology  4Google DeepMind
1 Introduction
Mathematical reasoning is fundamental to artificial intelligence (Newell and Simon, 1956), enabling significant progress in domains such as mathematics education, formally verified software, and mathematical research itself (Li et al., 2024). Recent advancements show that LLMs, when guided by careful prompting and tool integration, excel in solving straightforward high-school mathematics problems, but substantial challenges remain in problems requiring formal proof or the construction and verification of complex mathematical objects (Yang et al., 2024). MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer constructions.
Problems in math competitions broadly fall into two categories: theorem-proving and answer-construction tasks. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
Answer-Construction Problem:
Find all functions $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ such that
$\forall x,y\in\mathbb{R}^{+}$, $x\bigl{(}f(x)+f(y)\bigr{)}\geq\bigl{(}f(f(x))+y\bigr{)}f(y)$ Answer: All $f(x)=\frac{c}{x}$ for some $c>0$. Proof:
Theorem-Proving Problem:
Let $x_{1},\dots,x_{2023}$ be distinct positive numbers such that for all $n\in\{1,\dots,2023\}$,
$\sqrt{\left(\frac{1}{x_{1}}+\dots+\frac{1}{x_{n}}\right)\,(x_{1}+\dots+x_{n})}\in\mathbb{Z}$ Prove that $x_{2023}\geq 2024$. Proof:
Table 1: Examples of competitional-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021). While LLMs are effective at creative and exploratory reasoning, they frequently generate unverifiable and incorrect assertions due to hallucinations (Huang et al., 2025). Symbolic reasoning tools, such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008)), serve as the logical backbone of automated reasoning tools but cannot feasibly discover closed-form answers in the vast space of candidates.
Inspired by Mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer construction problems in math competition; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem proving techniques.
2 Related Work
Formal Theorem Proving
Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logic reasoning. Interactive theorem provers (ITPs) like Isabelle (Paulson, 1994) and Lean (Moura and Ullrich, 2021) allow step-by-step formal proof development, often integrating automation via tools such as Sledgehammer (Böhme and Nipkow, 2010) and LeanCopilot (Song et al., 2024).
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench. Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
LLMs for Mathematical Reasoning
Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs. Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity. Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems. Reinforcement learning-based approaches  (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals. Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools. SATLM (Ye et al., 2023) extends this framework by using an Z3 symbolic solver to validate logic-based task specifications, though its primary focus is on formal logic tasks rather than mathematics.
Evaluation Benchmarks
Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format. Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision. In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation. Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light. LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalize the Mathlib4 corpus back to natural language. PutnamBench remains actively maintained for recent versions of Lean.
3 Methods
3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove $P$ implies $Q$. Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
In contrast to theorem-proving tasks, an answer-construction task asks for constructing mathematical objects that satisfy specified constraints, without an explicitly stated conclusion. We define an answer-construction task by the following components:
Context Variables:
The unassigned context variables $a:=(a_{1},\dots,a_{k})$ (with type $\alpha:=\alpha_{1}\times\dots\times\alpha_{k}$) in the problem.
Context Property:
The predicate $P:\alpha\rightarrow\text{Prop}$ that context variables satisfy.
Answer Variables:
The unknown answer $b$ (with type $\beta$) to be constructed.
Answer Function:
$f:\alpha\to\text{Set }\beta$ mapping each context instance to a set of valid answers.
Problem Predicate:
a Boolean predicate $Q:\alpha\times\beta\to\text{Prop}$ such that $P(a,b)=\text{True}$ if and only if $b$ is a valid answer under context $a$.
We demonstrate the definition by a simple straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0,$ find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables will be $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$, the answer variable is $x$ with type $\mathbb{R}$, problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated as below in Lean’s type dependence theory:
$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a)$
[…]
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in closed-form and algebraically simplified representation that generalizes across instances. Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g. setting $f(a,b)=\{x|ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}|x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1. Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover. We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification. Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
3.2 Autoformalization Pipeline
Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean. Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail at precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, illustrated in Figure 1. The pipeline comprises three core components: (1) Autoformalizations with LLM (2) Lean knowledge retrieval and (3) error-guided feedback loops, and (3)
Figure 1: Overview of the autoformalization pipeline with error-based feedback and documentation retrieval: An informal statement is first processed by the LLM to produce a formalization. If the result fails either the Lean compilation check or the semantic check by an LLM judge, error messages and relevant Lean documentation are retrieved and fed back to the LLM. The process continues iteratively until the LLM produces a formalization that passes both checks.
Autoformalizations with LLM.
We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini, and deepseek-v3, for autoformalization. To better align these models with formalizations in high-school and undergraduate mathematics , we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023). Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
Lean Knowledge Retrieval.
LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated training data or scarce training data in formalized statements. To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023). After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries. Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024). When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances the formalization accuracy and reduces hallucination rates.
Error-Guided Feedback Loop.
Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies. Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations. This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers. Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, which ensures consistency across the formalization process. Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
3.3 Enumerate–Conjecture–Prove Framework
ECP systematically integrates LLM’s capabilities of coding, mathematical reasoning with formal theorem proving within Lean proof assistant. Figure 2 visualizes the workflow of ECP applied to an example problem.
Figure 2: Illustration of the ECP framework applied to a Balkan MO Shortlist problem in ConstructiveBench.
Enumerate Stage.
In the initial Enumerate stage, the “enumerator" LLM generates executable Python scripts designed to enumerate explicit examples or special cases of the given mathematical conditions, using the prompts in Appendix B. These enumeration scripts systematically evaluate finite subsets of context variables under computational constraints (timeout of 60 seconds and a maximum of 100 candidate answers) to gather concrete data points and identify initial numerical or structural patterns.
Conjecture Stage.
Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern finding in enumerated answers. These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification. We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
Prove Stage.
In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task. The theorem is then verified using specialized LLM fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025). Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
  1. 1.
    simp: Simplifies expressions by rewriting
  2. 2.
    aesop (Limperg and From, 2023): Proves goal via best-first search proof engine.
  3. 3.
    nlinarith: Solves nonlinear arithmetic problems over the integers and rationals.
  4. 4.
    ring: Proves equalities in commutative (semi)-rings.
  5. 5.
    norm_num: Evaluates and proves numeral equalities and inequalities.
Successful verification by Lean concludes the ECP process.
4 ConstructiveBench Dataset
In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination.
4.1 Data Sources and Scope
ConstructiveBench is curated by aggregating and rigorously filtering problems from reputable mathematical sources. Primary sources include official competition archives such as AMC 12 A/B (from 2000 onward), AIME I/II, HMMT, regional olympiads, and the IMO Shortlist and Longlist. Additionally, we integrate data from established informal math datasets including OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024).
The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions. It covers diverse mathematical domains including Algebra, Number Theory, Combinatorics, Probability, and Calculus. Rather than explicit domain labeling, we employ structural filtering to exclusively retain answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
((a)) Problem sources
((b)) Problem domains
Figure 3: Problem sources and problem domains in ConstructiveBench. Only the top 11 problem sources and top 7 problem domains are shown; the remaining are grouped under ‘Other’.
4.2 Data Deduplication
The dataset initially aggregated 5,555 problems. To remove duplication, both string-based matching and embedding-based duplicate filtering methods were utilized, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold. This process resulted in 5,316 unique and diverse problems.
4.3 Problem Autoformalization
We utilized our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into formally stated Lean theorems. Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with gpt-4o model serving as an LLM judge to verify semantic consistency.
Each problem went up to $T=5$ refinement iterations per model. This procedure yielded a well-formed subset of 3,431 problems, whose formalizations are compilable in the Lean compiler, passed LLM-based semantic check, and formatted according to the template. Each entry of ConstructiveBench includes the informal and formal problem statement and ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty. Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench. The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various mathematical types of problems.
Furthermore, to address potential data contamination, where LLMs might inadvertently have memorized training examples, we explicitly include a carefully selected test subset of 92 problems from the formalized data created after June 2024, beyond the knowledge cutoff date of most LLMs, including the ones that will be evaluated (gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
Figure 4: Left: Distribution of answer types in ConstructiveBench. Only the top 8 types are shown; the remaining types are grouped under ‘Other’. Right: Example dataset entry from ConstructiveBench.
5 Empirical Evaluation
The objective of our empirical evaluation was to answer the following questions.
  1. 1.
    Autoformalization Quality: Can ECP efficiently produce error-free and semantically faithful formalizations?
  2. 2.
    Answer-Construction Accuracy: Can ECP enable more accurate answer construction compared to CoT baselines?
  3. 3.
    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success on answer-construction problems compared to purely symbolic or LLM-only approaches?
5.1 Experimental Setup
Autoformalization
To evaluate the quality and efficiency of our autoformalization pipeline, we measure the percentage of problems that successfully compile in Lean and are validated by our LLM judge (gpt-4o). We sequentially run up to $T=5$ iterations for three models: fine-tuned gpt-4o-mini, deepseek-math-7b-finetune, and deepseek-v3.
Answer-Construction
For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically-correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers). We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, and deepseek-v3. The models are given the formalized lean statement, and the type of answer to be constructed. We compare the ECP framework against baseline Chain-of-Thought (CoT) method (i.e. without enumerating and pattern finding) for answer construction. The constructed answer is considered correct iff the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics simp, aesop, nlinarith, ring, norm_num).
Theorem-Proving
We evaluate these theorem-proving tasks using state-of-the-art theorem provers, comparing baseline symbolic methods and specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B. Evaluations are performed with one A100 GPU and 64 CPU cores for Lean verification, using hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048 and Pass@@@@32 metric. The experiment includes 2 conjecturing setups (CoT or ECP), 3 different models for conjecturing, and 3 different prover models for proving, resulting in a total of 18 experiments. We use 64 AMD EPYC 7453 processor CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours using an A100 80GB GPU for 2 hours per setup.
5.2 Results
Autoformalization
Out of the 5,316 problems in the ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks and, therefore, yielding an overall success rate of 64.6%. The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
Answer-Construction Accuracy
We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified problems from the ConstructiveBench and the subset of 337 problems that require answer construction from 640 problems in PutnamBench. Answer correctness was rigorously verified through formal proofs in Lean.
The results, summarized in Table 2, clearly demonstrate ECP’s advantage over standard LLM-only methods. On the ConstructiveBench dataset, ECP substantially improved answer accuracy for gpt-4.1-mini from 14.54% (baseline) to 45.06%, gpt-4.1-nano from 17.02% to 32.59%, and deepseek-v3 from 19.06% to 43.81%. Notably, gpt-4.1-mini achieved the highest accuracy within the ECP framework.
On the uncontaminated subset of 92 problems created after June 2024, performance improvements were also pronounced: gpt-4.1-mini improved from 5.43% to 31.52%, gpt-4.1-nano from 5.43% to 20.65%, and deepseek-v3 from 4.35% to 29.35%. These significant gains demonstrate ECP’s robust generalization capabilities, particularly for novel problems beyond the training cutoff.
Model / Dataset ConstructiveBench ConstructiveBench (After Cutoff) PutnamBench
gpt-4.1-nano 17.02% / 32.59% 5.43% / 20.65% 7.72% / 11.57%
gpt-4.1-mini 14.54% / 45.06% 5.43% / 31.52% 15.13% / 23.15%
deepseek-v3 19.06% / 43.81% 4.35% / 29.35% 18.99% / 31.75%
Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and 337 PutnamBench answer-construction problems between CoT baseline and ECP-assisted methods.
5.3 Theorem-Proving Accuracies
We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem-proving using the same ConstructiveBench problems. Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
Prover model
Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimini-Prover-7B-Preview DeepSeek-Prover-V2-7B
gpt-4.1-nano CoT 7.41% 10.46% 11.45% 15.42%
ECP 8.11% 13.61% 16.03% 20.61%
gpt-4.1-mini CoT 6.42% 8.07% 9.24% 13.00%
ECP 9.86% 15.83% 18.51% 25.01%
deepseek-v3 CoT 6.04% 9.62% 10.49% 15.33%
ECP 9.13% 20.26% 18.27% 24.63%
Table 3: Theorem proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
The summarized results on the ConstructiveBench problems are presented in Table 3. Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini. Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B. Additionally, on the subset of 337 problems requiring answer construction from PutnamBench, by using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, outperforming to 0.59% accuracy with CoT for answer construction. Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model-prover combinations.
6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. Our framework and dataset depend heavily on the quality of autoformalization produced by black-box LLMs, which are often trained on limited aligned informal-formal mathematical data (Wu et al., 2022b). In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity, it is time-consuming and difficult to scale. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence to the original problem. Future work should explore improved evaluation metrics for formalization quality, as well as expanding datasets that better align informal and formal representations of mathematical problems.
Theorem-proving: Our proving stage uses the off-the-shelf theorem provers, which leaves significant room for optimization. In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could lead to improvements in success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phases for difficult IMO problems, where the time complexity of the enumeration program blows up as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space. Fine-tuning LLMs or developing search-based algorithms on the formal answer construction task thus represents a promising avenue for next-generation mathematical AI models.
7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedbacks. In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%. Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer construction mathematics problems. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.
Appendix
Appendix A Lean Definition Filtering
To filter out the rarely-used definitions in autoformalization, we prioritize the definitions and theorems within the following namespaces in Lean:
Nat, Int, Rat, Real, Complex, ENat, NNReal, EReal, Monoid, CommMonoid,
Group, CommGroup, Ring, CommRing, Field, Algebra, Module, Set,  Finset,
Fintype, Multiset, List, Fin, BigOperators, Filter, Polynomial, Order,
SimpleGraph, Equiv, Embedding, Injective, Surjective, Bijective, Topology
Appendix B Prompt
B.1 Prompt for Formalization
[…]
‘‘‘lean
your code
B.2 Prompt for Enumerator Coding
LLM Prompt I have a difficult high-school competition-level math problem. Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer. Use the following format for the delimited output:
‘‘‘python
your code
B.3 Prompt for Conjecturing
LLM Prompt I have a difficult high-school competition-level math problem in Lean, which asks to find a answer set with a mathematical proof. Your task is to clearly state the final answer in Lean exprssion with explanations.You do not have to derive the whole proof. I have enumerated the answer(s) via program as a hint for you, and you may look at the pattern in it to help propose the final answer. Your final answer should be in closed form in Lean and generalized from enumerated answer that can eliminate the existential quantifier. You must not echo the problem statement in your answer. Use the following format for the delimited output:
‘‘‘lean
your code

### v3

@@
 Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery.
-Mathematical competitions, in particular, presents two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects.
+Mathematical competitions, in particular, present two challenging problem types: theorem-proving, which requires rigorous proofs of stated conclusions, and answer-construction, which involves hypothesizing and formally verifying mathematical objects.
 Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation.
@@
 We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations.
-On ConstructiveBench dataset, ECP improves the accuracy of the answer construction from Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with gpt-4.1-mini model.
-Moreover, combining with ECP’s constructed answers, state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines.
+On the ConstructiveBench dataset, ECP improves the accuracy of answer construction from a Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with the gpt-4.1-mini model.
+Moreover, combined with ECP’s constructed answers, the state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines.
 Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
@@
 MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer constructions.
-Problems in math competitions broadly fall into two categories: theorem-proving and answer-construction tasks.
+Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects.
 In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses.
@@
 $\sqrt{\left(\frac{1}{x_{1}}+\dots+\frac{1}{x_{n}}\right)\,(x_{1}+\dots+x_{n})}\in\mathbb{Z}$ Prove that $x_{2023}\geq 2024$. Proof:
-Table 1: Examples of competitional-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
+Table 1: Examples of competition-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
 We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021).
@@
 Inspired by mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework.
-Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem proving methods that verify these conjectures.
+Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures.
 ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
-Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer construction problems in math competition; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem proving techniques.
+Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer-construction problems in math competitions; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
 2 Related Work
 Formal Theorem Proving
-Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logic reasoning.
+Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logical reasoning.
 Interactive theorem provers (ITPs) like Isabelle (Paulson, 1994) and Lean (Moura and Ullrich, 2021) allow step-by-step formal proof development, often integrating automation via tools such as Sledgehammer (Böhme and Nipkow, 2010) and LeanCopilot (Song et al., 2024).
 Recent neuro-symbolic systems integrate LLMs with formal proof environments.
-Goedel-Prover (Lin et al., 2025), DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench.
+Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench.
 Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
@@
 Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools.
-SATLM (Ye et al., 2023) extends this framework by using an Z3 symbolic solver to validate logic-based task specifications, though its primary focus is on formal logic tasks rather than mathematics.
+SATLM (Ye et al., 2023) extends this framework by using a Z3 symbolic solver to validate logic-based task specifications.
 Evaluation Benchmarks
@@
 Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light.
-LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalize the Mathlib4 corpus back to natural language.
-PutnamBench remains actively maintained for recent versions of Lean.
+LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalizes the Mathlib4 corpus back to natural language.
 3 Methods
@@
 Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks.
-In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove $P$ implies $Q$.
+In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$.
 Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
@@
 Context Property:
-The predicate $P:\alpha\rightarrow\text{Prop}$ that context variables satisfy.
+The predicate $P:\alpha\rightarrow\text{Prop}$ that the context variables satisfy.
 Answer Variables:
@@
 A Boolean predicate $Q:\alpha\times\beta\to\text{Prop}$ such that $P(a,b)=\text{True}$ if and only if $b$ is a valid answer under context $a$.
-We demonstrate the definition by a simple straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0,$ find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables will be $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$, the answer variable is $x$ with type $\mathbb{R}$, problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
-In general, the task of answer-construction can be formulated as below in Lean’s type dependence theory:
-$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a)$
+We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$.
+The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
+In general, the task of answer-construction can be formulated as follows in Lean’s dependent type theory:
+$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a).$
 […]
-To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in closed-form and algebraically simplified representation that generalizes across instances.
-Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g.
-setting $f(a,b)=\{x|ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}|x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1.
+$\exists f:\alpha\rightarrow\beta,\ \forall(a:\alpha),\ P(a)\rightarrow\bigl{(}\forall(b:\beta),\,Q(a,b)\leftrightarrow b=f(a)\bigr{)}.$
+When there are no context variables, the formulation reduces to $\exists f:\text{Set }\beta,\ \forall b:\beta,\ P(b)\leftrightarrow b\in f.$
+To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in a closed-form, algebraically simplified representation that generalizes across instances.
+Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g., setting $f(a,b)=\{x\mid ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}\mid x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1.
 Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover.
@@
 Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean.
-Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail at precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
-To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, illustrated in Figure 1.
-The pipeline comprises three core components: (1) Autoformalizations with LLM (2) Lean knowledge retrieval and (3) error-guided feedback loops, and (3)
+Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail to produce precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
+To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, as illustrated in Figure 1.
+The pipeline comprises three core components: (1) autoformalizations with LLMs, (2) Lean knowledge retrieval, and (3) error-guided feedback loops.
 Figure 1: Overview of the autoformalization pipeline with error-based feedback and documentation retrieval: An informal statement is first processed by the LLM to produce a formalization.
@@
 Autoformalizations with LLM.
-We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini, and deepseek-v3, for autoformalization.
+We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini and deepseek-v3, for autoformalization.
 To better align these models with formalizations in high-school and undergraduate mathematics, we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023).
@@
 Lean Knowledge Retrieval.
-LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated training data or scarce training data in formalized statements.
-To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023).
+LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated or scarce training data in formalized statements.
+To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and the Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023).
 After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries.
 Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024).
-When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances the formalization accuracy and reduces hallucination rates.
+When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances formalization accuracy and reduces hallucination rates.
 Error-Guided Feedback Loop.
@@
 This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers.
-Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, which ensures consistency across the formalization process.
+Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, ensuring consistency across the formalization process.
 Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
 3.3 Enumerate–Conjecture–Prove Framework
-ECP systematically integrates LLM’s capabilities of coding, mathematical reasoning with formal theorem proving within Lean proof assistant.
+ECP systematically integrates LLMs’ capabilities for coding and mathematical reasoning with formal theorem proving within the Lean proof assistant.
 Figure 2 visualizes the workflow of ECP applied to an example problem.
@@
 Conjecture Stage.
-Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern finding in enumerated answers.
+Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern-finding in enumerated answers.
 These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification.
@@
 In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task.
-The theorem is then verified using specialized LLM fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025).
+The theorem is then verified using specialized LLMs fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025).
 Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
   1. 1.
-    simp: Simplifies expressions by rewriting
+    simp: Simplifies expressions by rewriting.
   2. 2.
-    aesop (Limperg and From, 2023): Proves goal via best-first search proof engine.
+    aesop (Limperg and From, 2023): Proves the goal via a best-first search proof engine.
   3. 3.
@@
 It covers diverse mathematical domains including algebra, number theory, combinatorics, probability, and calculus.
-Rather than explicit domain labeling, we employ structural filtering to exclusively retain answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
+Rather than explicit domain labeling, we employ structural filtering to retain exclusively answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
 ((a)) Problem sources
@@
 The dataset initially aggregated 5,555 problems.
-To remove duplication, both string-based matching and embedding-based duplicate filtering methods were utilized, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold.
+To remove duplication, we utilized both string-based matching and embedding-based duplicate filtering methods, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold.
 This process resulted in 5,316 unique and diverse problems.
 4.3 Problem Autoformalization
-We utilized our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into formally stated Lean theorems.
-Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with gpt-4o model serving as an LLM judge to verify semantic consistency.
-Each problem went up to $T=5$ refinement iterations per model.
-This procedure yielded a well-formed subset of 3,431 problems, whose formalizations are compilable in the Lean compiler, passed LLM-based semantic check, and formatted according to the template.
-Each entry of ConstructiveBench includes the informal and formal problem statement and ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty.
+We used our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into a formally stated Lean theorem.
+Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with the gpt-4o model serving as an LLM judge to verify semantic consistency.
+Each problem underwent up to $T=5$ refinement iterations per model.
+This procedure yielded a well-formed subset of 3,431 problems, whose formalizations compile successfully in Lean, pass the LLM-based semantic check, and are formatted according to the template.
+Each entry in ConstructiveBench includes the informal and formal problem statements, the ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty.
 Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench.
-The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various mathematical types of problems.
-Furthermore, to address potential data contamination, where LLMs might inadvertently have memorized training examples, we explicitly include a carefully selected test subset of 92 problems from the formalized data created after June 2024, beyond the knowledge cutoff date of most LLMs, including the ones that will be evaluated (gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
+The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various types of mathematical problems.
+Furthermore, to address potential data contamination—where LLMs might inadvertently have memorized training examples—we explicitly include a carefully selected test subset of 92 problems drawn from formalized data created after June 2024, which is beyond the knowledge cutoff date of most LLMs (including those we will evaluate: gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
 Figure 4: Left: Distribution of answer types in ConstructiveBench. Only the top 8 types are shown; the remaining types are grouped under ‘Other’. Right: Example dataset entry from ConstructiveBench.
@@
   3. 3.
-    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success on answer-construction problems compared to purely symbolic or LLM-only approaches?
+    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success rates on answer-construction problems compared to purely symbolic or LLM-only approaches?
 5.1 Experimental Setup
-Autoformalization
+Autoformalization.
 To evaluate the quality and efficiency of our autoformalization pipeline, we measure the percentage of problems that successfully compile in Lean and are validated by our LLM judge (gpt-4o).
 We sequentially run up to $T=5$ iterations for three models: fine-tuned gpt-4o-mini, deepseek-math-7b-finetune, and deepseek-v3.
-Answer-Construction
-For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically-correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers).
-We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, and deepseek-v3.
-The models are given the formalized lean statement, and the type of answer to be constructed.
-We compare the ECP framework against baseline Chain-of-Thought (CoT) method (i.e.
-without enumerating and pattern finding) for answer construction.
-The constructed answer is considered correct iff the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics simp, aesop, nlinarith, ring, norm_num).
-Theorem-Proving
-We evaluate these theorem-proving tasks using state-of-the-art theorem provers, comparing baseline symbolic methods and specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B.
-Evaluations are performed with one A100 GPU and 64 CPU cores for Lean verification, using hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048 and Pass@@@@32 metric.
-The experiment includes 2 conjecturing setups (CoT or ECP), 3 different models for conjecturing, and 3 different prover models for proving, resulting in a total of 18 experiments.
-We use 64 AMD EPYC 7453 processor CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours using an A100 80GB GPU for 2 hours per setup.
+Answer-Construction.
+For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers).
+We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, as well as deepseek-v3.
+The models are given the formalized Lean statement and the type of answer to be constructed.
+We compare the ECP framework against a baseline Chain-of-Thought (CoT) method (i.e., without enumeration and pattern finding) for answer construction.
+The constructed answer is considered correct if and only if the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics: simp, aesop, nlinarith, ring, norm_num).
+Theorem-Proving.
+We evaluate the theorem-proving tasks using state-of-the-art provers, comparing baseline symbolic methods with specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B.
+Evaluations are performed on one A100 GPU and 64 CPU cores for Lean verification, using the hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048, and the Pass@32 metric.
+The experiment includes two conjecturing setups (CoT or ECP), three different models for conjecturing, and three different prover models for proving, resulting in a total of 18 experiments.
+We use 64 AMD EPYC 7453 CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours on an A100 80 GB GPU (2 hours per setup).
 5.2 Results
-Autoformalization
-Out of the 5,316 problems in the ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks and, therefore, yielding an overall success rate of 64.6%.
+Autoformalization.
+Out of the 5,316 problems in ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks, yielding an overall success rate of 64.6%.
 The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
-Answer-Construction Accuracy
-We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified problems from the ConstructiveBench and the subset of 337 problems that require answer construction from 640 problems in PutnamBench.
+Answer-Construction.
+We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified ConstructiveBench problems and on the subset of 337 answer-construction problems from the 640 problems in PutnamBench.
 Answer correctness was rigorously verified through formal proofs in Lean.
@@
 deepseek-v3 19.06% / 43.81% 4.35% / 29.35% 18.99% / 31.75%
-Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and 337 PutnamBench answer-construction problems between CoT baseline and ECP-assisted methods.
-5.3 Theorem-Proving Accuracies
-We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem-proving using the same ConstructiveBench problems.
+Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and on 337 PutnamBench answer-construction problems.
+Theorem-Proving.
+We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem proving using the same ConstructiveBench problems.
 Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
 Prover Model
-Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimini-Prover-7B-Preview DeepSeek-Prover-V2-7B
+Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimina-Prover-7B-Preview DeepSeek-Prover-V2-7B
 gpt-4.1-nano CoT 7.41% 10.46% 11.45% 15.42%
@@
 ECP 9.13% 20.26% 18.27% 24.63%
-Table 3: Theorem proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
+Table 3: Theorem-proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
 The summarized results on the ConstructiveBench problems are presented in Table 3.
@@
 Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B.
-Additionally, on the subset of 337 problems requiring answer construction from PutnamBench, by using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, outperforming to 0.59% accuracy with CoT for answer construction.
-Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model-prover combinations.
+Additionally, on the subset of 337 answer-construction problems from PutnamBench, using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, compared to 0.59% with CoT for answer construction.
+Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model–prover combinations.
 6 Limitations and Future Work
 Quality of Autoformalized Theorem Statements.
-Our framework and dataset depend heavily on the quality of autoformalization produced by black-box LLMs, which are often trained on limited aligned informal-formal mathematical data (Wu et al., 2022b).
+Our framework and dataset depend on the quality of autoformalization produced by black-box LLMs.
 In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics.
 While manual evaluation offers high fidelity, it is time-consuming and difficult to scale.
-Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence to the original problem.
-Future work should explore improved evaluation metrics for formalization quality, as well as expanding datasets that better align informal and formal representations of mathematical problems.
-Theorem-proving: Our proving stage uses the off-the-shelf theorem provers, which leaves significant room for optimization.
-In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could lead to improvements in success rates and runtime.
-Another limitation is the combinatorial explosion during the enumeration phases for difficult IMO problems, where the time complexity of the enumeration program blows up as the in-context parameter increases.
+Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem.
+Future work should explore improved evaluation metrics for formalization quality and expand datasets that better align informal and formal representations of mathematical problems.
+Theorem-Proving.
+Our proving stage uses off-the-shelf theorem provers, which leaves significant room for optimization.
+In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime.
+Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases.
 Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space.
-Fine-tuning LLMs or developing search-based algorithms on the formal answer construction task thus represents a promising avenue for next-generation mathematical AI models.
+Fine-tuning LLMs or developing search-based algorithms for the formal answer-construction task thus represents a promising avenue for next-generation mathematical AI models.
 7 Conclusion
 We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean.
-To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedbacks.
+To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedback.
 In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%.
 Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end.
-These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer construction mathematics problems.
+These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer-construction mathematics problems.
 Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.

### v4

@@
 Abstract
-Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery.
-Mathematical competitions, in particular, present two challenging problem types: theorem-proving, which requires rigorous proofs of stated conclusions, and answer-construction, which involves hypothesizing and formally verifying mathematical objects.
-Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation.
-We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving.
-We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations.
-On the ConstructiveBench dataset, ECP improves the accuracy of answer construction from a Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with the gpt-4.1-mini model.
-Moreover, combined with ECP’s constructed answers, the state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines.
-Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
-1University of Toronto  2Vector Institute  3Georgia Institute of Technology  4Google DeepMind
+Mathematical reasoning is central to artificial intelligence, with applications in education, code generation, and research-level mathematical discovery.
+Mathematical competitions highlight two problem types: theorem-proving, requiring rigorous proofs, and answer-construction, requiring creative generation and formal verification of mathematical objects.
+Existing research reveals that LLMs can tackle difficult answer-construction tasks but are prone to errors from hallucinations and unverifiable steps, while symbolic methods guarantee rigor but falter in creative answer construction.
+This raises a key understudied question: how to solve answer-construction problems while preserving both LLM creativity and mathematical rigor?
+To address this problem, we introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving in Lean, and ConstructiveBench, a dataset of 3,640 formal answer-construction problems from math competitions.
+ECP is model-agnostic and shows consistent improvements over pure LLM baselines: on the subset of PutnamBench for answer construction, ECP formally solves 6 out of 337 answer-construction problems end-to-end (up from 4 without ECP) using GPT-5 mini and DeepSeek-Prover-V2-7B.
+On ConstructiveBench, ECP achieves 33.1% end-to-end state-of-the-art accuracy (up from 32.5%), demonstrating its potential to advance formal mathematical reasoning by combining LLM conjecturing with formal verification.
+Our code and dataset are publicly available at GitHub [^1] https://github.com/sunjia72/ECP and Hugging Face [^2] https://huggingface.co/datasets/sunjia72/ConstructiveBench.
+1University of Toronto    2Vector Institute    3Georgia Institute of Technology
 1 Introduction
@@
 Recent advancements show that LLMs, when guided by careful prompting and tool integration, excel in solving straightforward high-school mathematics problems, but substantial challenges remain in problems requiring formal proof or the construction and verification of complex mathematical objects (Yang et al., 2024).
-MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer constructions.
+MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer construction.
 Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects.
 In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses.
-Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness.
+Conversely, answer-construction problems require the construction of specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness.
 Representative examples of these problems appear in Table 1.
-Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
+Crucially, answer-construction problems present the unique challenge of identifying the correct candidate object or structure prior to formal verification.
 Answer-Construction Problem:
@@
 $\sqrt{\left(\frac{1}{x_{1}}+\dots+\frac{1}{x_{n}}\right)\,(x_{1}+\dots+x_{n})}\in\mathbb{Z}$ Prove that $x_{2023}\geq 2024$. Proof:
-Table 1: Examples of competition-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
+Table 1: Examples of competition-level problems. The left panel shows an answer-construction problem requiring a candidate answer and proof. The right panel shows a standard theorem-proving problem.
 We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021).
+Autoformalization is an automated process that translates informal problem statements into formal representations.
 While LLMs are effective at creative and exploratory reasoning, they frequently generate unverifiable and incorrect assertions due to hallucinations (Huang et al., 2025).
 Symbolic reasoning tools, such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008)), serve as the logical backbone of automated reasoning tools but cannot feasibly discover closed-form answers in the vast space of candidates.
-Inspired by mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework.
+Inspired by Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the ECP framework.
 Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures.
 ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
-Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer-construction problems in math competitions; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
+Our contributions are twofold: (i) ECP, a modular framework that unifies LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving in Lean; (ii) ConstructiveBench, an autoformalized dataset of 3640 competition-level answer-construction problems with verified Lean formalizations, validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
 2 Related Work
@@
 Recent neuro-symbolic systems integrate LLMs with formal proof environments.
-Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench.
-Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
+Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), InternLM-Math (Ying et al., 2024b), SeedProver (Chen et al., 2025), Kimina-Prover (Wang et al., 2025) and AlphaProof (Teams, 2024) scale formal data generation and RL-style training to improve Lean proof success on miniF2F, PutnamBench, and related evaluations.
+Domain-specific systems, including AlphaGeometry (Trinh et al., 2024; Chervonyi et al., 2025) and PyEuclid (Li et al., 2025b) for Euclidean geometry, LIPS for inequalities (Li et al., 2025a), target particular subfields with symbolic search and automation.
+Beyond theorem proving,  Liu et al.
+(2025b) introduce a search-based formal problem-solving framework (FPS/D-FPS) with new benchmarks for process-verified solutions.
 LLMs for Mathematical Reasoning
-Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs.
-Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity.
-Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems.
-Reinforcement learning-based approaches (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals.
-Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools.
-SATLM (Ye et al., 2023) extends this framework by using a Z3 symbolic solver to validate logic-based task specifications.
+Beyond formal proof generation, the broader math-reasoning literature explores prompting, learning, and tool use.
+Prompting-based approaches (e.g., chain-of-thought) encourage explicit intermediate reasoning (Wei et al., 2023).
+Reinforcement learning has recently become central: GRPO—popularized in Shao et al.
+(2024)—and large-scale systems like DAPO (Yu et al., 2025a) improve step-wise reasoning via preference/advantage signals at scale.
+Complementary to policy learning, tool-integrated multi-turn training frameworks such as ReTool (Feng et al., 2025) explicitly train agents to plan, call tools, and verify intermediate computations.
+Orthogonal “code-as-tool” frameworks, including PAL (Gao et al., 2023) and Tora (Gou et al., 2023), and logic-solver integration such as SATLM (Ye et al., 2023), demonstrate the benefits of executing verifiable intermediate steps.
+Liu et al.
+(2025b) propose a broader formal problem-solving framework (FPS/D-FPS) with new benchmarks for process-verified solutions.
 Evaluation Benchmarks
-Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format.
-Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision.
-In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation.
-Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light.
-LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalizes the Mathlib4 corpus back to natural language.
+We group evaluation datasets into informal and formal categories.
+Early works in mathematical evaluations, such as GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) assess numerical/short-answer problem solving in natural language.
+Follow-up works include harder answer-only suites like OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) though they still rely on string matching for answer verification.
+In parallel, formal benchmarks like miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), ProofNet (Azerbayev et al., 2023), LeanWorkBook (Ying et al., 2024a), Herald (Gao et al., 2024b), and the recent FormalMATH (Yu et al., 2025b) enable machine-verifiable evaluation in Lean/HOL.
+ProverBench (Xin et al., 2024) contributes a mixed set of 325 formalized problems, including AIME-style items, expanding coverage toward competition math.
+MathConstruct (Balunović et al., 2025) targets constructive proof tasks.
+CombiBench (Liu et al., 2025a) focuses on combinatorics in Lean.
 3 Methods
@@
 In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$.
-Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
-In contrast to theorem-proving tasks, an answer-construction task asks for constructing mathematical objects that satisfy specified constraints, without an explicitly stated conclusion.
-We define an answer-construction task by the following components:
+Prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this theorem-proving task format, which aligns with traditional first-order theorem proving.
+In contrast to theorem-proving tasks, an answer-construction task asks for the construction of mathematical objects satisfying specified constraints, without an explicitly stated conclusion.
+We define an answer-construction task by the following components in the context of Lean dependent type theory:
 Context Variables:
@@
 A Boolean predicate $Q:\alpha\times\beta\to\text{Prop}$ such that $P(a,b)=\text{True}$ if and only if $b$ is a valid answer under context $a$.
-We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$.
+We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all $x\in\mathbb{R}$ such that ax=bax=b." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=a\neq 0$.
 The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
-In general, the task of answer-construction can be formulated as follows in Lean’s dependent type theory:
+In general, the task of answer-construction can be formulated in Lean’s dependent type theory as follows:
 $\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a).$
-[…]
+Intuitively, the task is to construct a function $f$ such that for any context parameter $a$ that satisfies property $P$, $f(a)$ precisely captures all and only the valid answers.
+Variants of format constraints on answers can be incorporated by refining the problem predicate $Q$.
+For example, to enforce minimality, it is sufficient to define $Q_{\mathit{min}}(a,b):=P(a,b)\wedge\forall c<b,\;\neg P(a,c)$ or use Lean’s built-in notions such as IsLeast or IsGreatest.
+When the problem admits a single unique solution, $f$ degenerates to $\alpha\to\beta$, and it becomes sufficient to show
 $\exists f:\alpha\rightarrow\beta,\ \forall(a:\alpha),\ P(a)\rightarrow\bigl(\forall(b:\beta),\,Q(a,b)\leftrightarrow b=f(a)\bigr).$
 When there are no context variables, the formulation reduces to $\exists f:\text{Set }\beta,\ \forall b:\beta,\ P(b)\leftrightarrow b\in f.$
-To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in a closed-form, algebraically simplified representation that generalizes across instances.
-Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g., setting $f(a,b)=\{x\mid ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}\mid x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1.
-Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover.
-We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification.
-Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
-3.2 Autoformalization Pipeline
-Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean.
-Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail to produce precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
-To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, as illustrated in Figure 1.
-The pipeline comprises three core components: (1) autoformalizations with LLMs, (2) Lean knowledge retrieval, and (3) error-guided feedback loops.
-Figure 1: Overview of the autoformalization pipeline with error-based feedback and documentation retrieval: An informal statement is first processed by the LLM to produce a formalization.
-If the result fails either the Lean compilation check or the semantic check by an LLM judge, error messages and relevant Lean documentation are retrieved and fed back to the LLM.
-The process continues iteratively until the LLM produces a formalization that passes both checks.
-Autoformalizations with LLM.
-We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini and deepseek-v3, for autoformalization.
-To better align these models with formalizations in high-school and undergraduate mathematics, we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023).
-Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
-Lean Knowledge Retrieval.
-LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated or scarce training data in formalized statements.
-To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and the Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023).
-After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries.
-Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024).
-When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances formalization accuracy and reduces hallucination rates.
-Error-Guided Feedback Loop.
-Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies.
-Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations.
-This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers.
-Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, ensuring consistency across the formalization process.
-Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
+To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in closed-form canonical format without echoing the problem statement’s syntax or structure.
+For example, in the following task: theorem test: (x y : $\mathbb{N}$) (hpos : 0 < x $\land$ 0 < y) : x ^ 3 + y ^ 3 = x ^ 2 + 42 * x * y + y ^ 2 $\leftrightarrow$ (x, y) $\in$ answer
+[…]
+The formal answer-construction task is expressed as a theorem statement with a placeholder or an existentially quantified answer.
+The autoformalization process will be discussed in Section 4.3 in detail.
 3.2 Enumerate–Conjecture–Prove Framework
-ECP systematically integrates LLMs’ capabilities for coding and mathematical reasoning with formal theorem proving within the Lean proof assistant.
-Figure 2 visualizes the workflow of ECP applied to an example problem.
-Figure 2: Illustration of the ECP framework applied to a Balkan MO Shortlist problem in ConstructiveBench.
+ECP enables LLM to perform multi-turn tool calls with Python and Lean in the ReAct manner (Yao et al., 2023).
+The framework and structured output are implemented using the APPL prompt programming language (Dong et al., 2024), and we include the prompt template in Appendix B.
+Figure 1 visualizes the workflow of ECP applied to an example answer-construction problem end-to-end.
+Figure 1: Illustration of the ECP framework applied to a formalized Balkan MO Shortlist problem in OmniMath dataset (Gao et al., 2024a).
 Enumerate Stage.
-In the initial Enumerate stage, the “enumerator" LLM generates executable Python scripts designed to enumerate explicit examples or special cases of the given mathematical conditions, using the prompts in Appendix B.
-These enumeration scripts systematically evaluate finite subsets of context variables under computational constraints (timeout of 60 seconds and a maximum of 100 candidate answers) to gather concrete data points and identify initial numerical or structural patterns.
+In the initial Enumerate stage, the LLM is given a formal answer-construction task, and the goal is to enumerate the candidate answers by generating Python programs.
+The model performs multi-turn interaction to refine the program and produce enumeration: {<reasoning_1> <python_call_1> <python_response_1><reasoning_2>…}.
+For problems with unassigned context variables or infinitely many answers, the model is encouraged to explore small customized setups and return bounded enumerations, guided by the system prompt with constraints of timeout=60s and a maximum of 100 enumerated answers.
 Conjecture Stage.
-Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern-finding in enumerated answers.
-These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification.
-We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
+In the Conjecture (i.e.
+answer-construction) stage, given the answer-construction task and the enumerations as hint, the LLM reasons about the problem, generalizes beyond enumeration and attempts to propose a syntactically valid closed-form answer in Lean, following the workflow {<reasoning_1> <lean_call_1> <lean_response_1><reasoning_2> …}.
+To prevent potential answer “hacking" described in Section  3.1, we design the system prompt to explicitly disallow such illegal answers that lead to trivial or vacuous versions of the Lean theorem statement.
 Prove Stage.
-In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task.
-The theorem is then verified using specialized LLMs fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025).
-Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
+In the final Prove stage, the candidate answer is substituted into the theorem statement, reducing the answer-construction problem into a standard theorem-proving task.
+Our framework supports invoking theorem-proving models, such as DeepSeek-Prover-V2 (Ren et al., 2025), Kimina-Prover (Wang et al., 2025), and Goedel-Prover (Lin et al., 2025), to generate proofs of the theorem, and Lean verifies whether the proof closes the goal.
+Additionally, we implement pure symbolic methods as a sequence of automation tactics as follows, which will be used to verify equivalence between constructed and ground-truth answer in evaluation:
   1. 1.
@@
     norm_num: Evaluates and proves numeral equalities and inequalities.
-Successful verification by Lean concludes the ECP process.
+Successful Lean verification of the proof concludes the ECP pipeline.
 4 ConstructiveBench Dataset
-In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination.
+In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination and a human evaluation of autoformalization quality.
 4.1 Data Sources and Scope
@@
 Additionally, we integrate data from established informal math datasets including OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024).
-The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions.
-It covers diverse mathematical domains including algebra, number theory, combinatorics, probability, and calculus.
-Rather than explicit domain labeling, we employ structural filtering to retain exclusively answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
+We employ structural filtering to retain exclusively answer-construction problems—those requiring numerical or symbolic answers alongside proofs—and exclude purely theorem-proving tasks without explicit answer-construction components.
+4.2 Data Deduplication
+The raw source dataset contains 5,555 problems.
+To remove duplicates, we apply both string-based matching and embedding-based filtering using Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic-similarity threshold.
+This process yields 5,316 unique problems spanning a broad difficulty range—from challenging high-school problems to sophisticated Olympiad-style questions that cover diverse domains including algebra, number theory, combinatorics, probability, and calculus.
 ((a)) Problem sources
 ((b)) Problem domains
-Figure 3: Problem sources and problem domains in ConstructiveBench. Only the top 11 problem sources and top 7 problem domains are shown; the remaining are grouped under ‘Other’.
-4.2 Data Deduplication
-The dataset initially aggregated 5,555 problems.
-To remove duplication, we utilized both string-based matching and embedding-based duplicate filtering methods, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold.
-This process resulted in 5,316 unique and diverse problems.
+Figure 2: Problem sources and domains in ConstructiveBench. Only the top 11 sources and top 7 domains are shown; the remainder are grouped under ‘Other’.
 4.3 Problem Autoformalization
-We used our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into a formally stated Lean theorem.
-Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with the gpt-4o model serving as an LLM judge to verify semantic consistency.
-Each problem underwent up to $T=5$ refinement iterations per model.
-This procedure yielded a well-formed subset of 3,431 problems, whose formalizations compile successfully in Lean, pass the LLM-based semantic check, and are formatted according to the template.
-Each entry in ConstructiveBench includes the informal and formal problem statements, the ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty.
-Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench.
-The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various types of mathematical problems.
-Furthermore, to address potential data contamination—where LLMs might inadvertently have memorized training examples—we explicitly include a carefully selected test subset of 92 problems drawn from formalized data created after June 2024, which is beyond the knowledge cutoff date of most LLMs (including those we will evaluate: gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
-Figure 4: Left: Distribution of answer types in ConstructiveBench. Only the top 8 types are shown; the remaining types are grouped under ‘Other’. Right: Example dataset entry from ConstructiveBench.
+To formally solve answer-construction problems in Lean, we first autoformalize each informal problem into the formal task described in Section 3.1.
+Autoformalization with LLMs.
+We leverage state-of-the-art autoformalization models Herald-Translator (Gao et al., 2024b) and Kimina-Autoformalizer-7B (Wang et al., 2025), together with the commercial model GPT-5 mini using a 3-shot prompt.
+We also use GPT-5 mini as an LLM judge.
+A candidate formalization is accepted if and only if both conditions hold: (i) the Lean compiler returns no errors; and (ii) the LLM judge deems the formalization semantically equivalent to the informal statement, with correct use of key definitions and no trivial restatement.
+Information Retrieval in Lean.
+To mitigate hallucinations and errors arising from scarce aligned informal-to-formal data (Wu et al., 2022) and rapidly evolving Lean syntax, we construct an external Lean knowledge base and an interactive loop for interaction between LLM and Lean.
+The knowledge base contains approximately 385K entries from the Lean 4.23.0 documentation and the mathlib repository (mathlib Community, 2020).
+After filtering to retain entries most relevant to high-school-level formalizations (Appendix A), the database is reduced to 57K items.
+Each entry is embedded with Sentence-BERT (Reimers and Gurevych, 2019) and indexed by FAISS (Douze et al., 2024); when error messages indicate missing or hallucinated definitions, we retrieve the top 5 candidates by semantic similarity and the top 5 by edit distance on the symbol name.
+Error-Guided Feedback Loop.
+Each of the three LLMs receives the informal problem with a 3-shot prompt and drafts a formalization.
+Lean compiler error messages, retrieved definitions, and LLM-judge feedback are then fed back to the model for iterative refinement, up to $T=5$ iterations.
+A problem is considered successfully autoformalized if any model produces an accepted formalization within the budget (a best-of-$N$ strategy across models).
+Figure 3: Overview of the autoformalization pipeline: the LLM drafts a formal statement; when compilation or semantic checks fail, error messages and retrieved Lean references are fed back; the loop repeats until both checks pass.
+4.4 Dataset Validation
+Outputs and Coverage.
+Starting from 5,316 deduplicated informal problems, we run up to $T=5$ refinement iterations per model.
+This produced 4,729 Lean-compilable formalizations in the correct format; among these, 3,640 also pass the LLM-judge semantic check and constituted the final ConstructiveBench dataset.
+Each accepted entry includes the informal statement, the formal Lean theorem, the ground-truth answer, an informal solution, and metadata such as source and difficulty.
+Figure 4 shows an example dataset entry and the distribution of answer types.
+The diversity of answer types enables comprehensive evaluation across a wide range of problem formats.
+Figure 4: Left: Distribution of answer types in ConstructiveBench (top 8 shown; the remainder are grouped as ‘Other’). Right: Example dataset entry.
+After-Cutoff Subset.
+To mitigate contamination from pretraining corpora—where LLMs might inadvertently memorize training examples—we include a test split of 106 problems created after June 2024 (beyond the knowledge cutoff date of most LLMs we will evaluate), ensuring fair evaluation on unseen problems.
+Human Evaluation.
+To estimate the quality of autoformalized theorem statements in ConstructiveBench, we conducted a human evaluation on a random slice of 100 formalizations from the 3,640 accepted problems.
+Following the rubric in Gao et al.
+(2024b), each formalization was labeled as completely correct, minor error, or major error.
+We found 77 were fully correct, 6 had minor errors, and 17 had major errors; examples and analyses appear in Appendix C.
+This result aligns with human evaluations of existing autoformalized benchmarks (66.9% accuracy on Herald (Gao et al., 2024b) and 72.1% on FormalMath (Yu et al., 2025b)), highlighting the high quality of ConstructiveBench.
 5 Empirical Evaluation
-The objective of our empirical evaluation was to answer the following questions.
+We evaluate whether ECP outperforms a classic chain-of-thought (CoT) baseline on two tasks:
   1. 1.
-    Autoformalization Quality: Can ECP efficiently produce error-free and semantically faithful formalizations?
+    Answer-Construction. Given a formal answer-construction problem, can the conjecturer produce a Lean expression provably equivalent to the ground truth under a fixed set of automation tactics?
   2. 2.
-    Answer-Construction Accuracy: Can ECP enable more accurate answer construction compared to CoT baselines?
-  3. 3.
-    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success rates on answer-construction problems compared to purely symbolic or LLM-only approaches?
+    End-to-End Problem Solving. Given a formal answer-construction problem, can the system both conjecture a correct answer and close the Lean goal using a prover?
 5.1 Experimental Setup
-Autoformalization.
-To evaluate the quality and efficiency of our autoformalization pipeline, we measure the percentage of problems that successfully compile in Lean and are validated by our LLM judge (gpt-4o).
-We sequentially run up to $T=5$ iterations for three models: fine-tuned gpt-4o-mini, deepseek-math-7b-finetune, and deepseek-v3.
-Answer-Construction.
-For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers).
-We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, as well as deepseek-v3.
-The models are given the formalized Lean statement and the type of answer to be constructed.
-We compare the ECP framework against a baseline Chain-of-Thought (CoT) method (i.e., without enumeration and pattern finding) for answer construction.
-The constructed answer is considered correct if and only if the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics: simp, aesop, nlinarith, ring, norm_num).
-Theorem-Proving.
-We evaluate the theorem-proving tasks using state-of-the-art provers, comparing baseline symbolic methods with specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B.
-Evaluations are performed on one A100 GPU and 64 CPU cores for Lean verification, using the hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048, and the Pass@32 metric.
-The experiment includes two conjecturing setups (CoT or ECP), three different models for conjecturing, and three different prover models for proving, resulting in a total of 18 experiments.
-We use 64 AMD EPYC 7453 CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours on an A100 80 GB GPU (2 hours per setup).
-5.2 Results
-Autoformalization.
-Out of the 5,316 problems in ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks, yielding an overall success rate of 64.6%.
-The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
-Answer-Construction.
-We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified ConstructiveBench problems and on the subset of 337 answer-construction problems from the 640 problems in PutnamBench.
-Answer correctness was rigorously verified through formal proofs in Lean.
-The results, summarized in Table 2, clearly demonstrate ECP’s advantage over standard LLM-only methods.
-On the ConstructiveBench dataset, ECP substantially improved answer accuracy for gpt-4.1-mini from 14.54% (baseline) to 45.06%, gpt-4.1-nano from 17.02% to 32.59%, and deepseek-v3 from 19.06% to 43.81%.
-Notably, gpt-4.1-mini achieved the highest accuracy within the ECP framework.
-On the uncontaminated subset of 92 problems created after June 2024, performance improvements were also pronounced: gpt-4.1-mini improved from 5.43% to 31.52%, gpt-4.1-nano from 5.43% to 20.65%, and deepseek-v3 from 4.35% to 29.35%.
-These significant gains demonstrate ECP’s robust generalization capabilities, particularly for novel problems beyond the training cutoff.
-Model / Dataset ConstructiveBench ConstructiveBench (After Cutoff) PutnamBench
-gpt-4.1-nano 17.02% / 32.59% 5.43% / 20.65% 7.72% / 11.57%
-gpt-4.1-mini 14.54% / 45.06% 5.43% / 31.52% 15.13% / 23.15%
-deepseek-v3 19.06% / 43.81% 4.35% / 29.35% 18.99% / 31.75%
-Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and on 337 PutnamBench answer-construction problems.
-Theorem-Proving.
-We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem proving using the same ConstructiveBench problems.
-Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
-Prover Model
-Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimina-Prover-7B-Preview DeepSeek-Prover-V2-7B
-gpt-4.1-nano CoT 7.41% 10.46% 11.45% 15.42%
-ECP 8.11% 13.61% 16.03% 20.61%
-gpt-4.1-mini CoT 6.42% 8.07% 9.24% 13.00%
-ECP 9.86% 15.83% 18.51% 25.01%
-deepseek-v3 CoT 6.04% 9.62% 10.49% 15.33%
-ECP 9.13% 20.26% 18.27% 24.63%
-Table 3: Theorem-proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
-The summarized results on the ConstructiveBench problems are presented in Table 3.
-Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini.
-Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B.
-Additionally, on the subset of 337 answer-construction problems from PutnamBench, using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, compared to 0.59% with CoT for answer construction.
-Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model–prover combinations.
+Datasets.
+We evaluate on ConstructiveBench (3,640 Lean-verified answer-construction problems) and the PutnamBench answer-construction subset (337 problems).
+We additionally report a contamination-controlled After-Cutoff split of 106 ConstructiveBench problems (created after June 2024).
+Models and Hyperparameters.
+For conjecturing we use GPT-5 (mini, nano), GPT-4.1 (mini, nano), and DeepSeek (V3.1, V3).
+For reasoning models (GPT-5 families and DeepSeek-V3.1) and lightweight non-reasoning models (GPT-4.1 families and DeepSeek-V3), we set reasoning_effort to medium and None, respectively.
+We use each model’s recommended decoding default parameters; for sampling we set max_tokens=4096, coder_max_attempt=3, and conjecturing_attempt=5.
+For proving, we use DeepSeek-Prover-V2-7B, Goedel-Prover-SFT, and Kimina-Prover-Preview-7B, with hyperparameters top_p=0.95, temperature=1.0, and max_tokens=4096.
+We evaluate using Pass@32 with a 120s Lean verification timeout.
+Metrics.
+Answer-construction accuracy marks a problem correct if the conjectured answer $a$ is provably equal to the ground truth $a_{\text{gt}}$ via the following Lean automation tactics: simp, aesop (Limperg and From, 2023), nlinarith, ring, and norm_num.
+End-to-end accuracy marks a problem correct if the answer is constructed by the conjecturer and the prover produces a Lean-verified proof to the theorem statement after substituting the conjectured answer.
+We use Pass@32 for prover, meaning that each prover model is sampled 32 outputs independently and is marked correct if at least one proof passes Lean check.
+Given the fixed model choice, union accuracy simply measures the fraction of successfully solved problems by any of CoT and ECP.
+5.2 Answer-Construction Accuracy
+In this section, we analyze the results of formal answer-construction accuracies with different conjecturer models on ConstructiveBench and PutnamBench shown in Table 2 and Table 3.
+Complete results for six models, including evaluations on the after-knowledge-cutoff dataset, are listed in Appendix D.
+Results show that ECP consistently improves answer construction across conjecturers and datasets.
+On ConstructiveBench, GPT-5 mini rises from 69.7% to 73.6%, and DeepSeek-V3.1 increases from 65.4% to 75.0%, indicating that stronger conjecturers still benefit from enumeration signals.
+The largest absolute gains occur for lightweight non-reasoning models: DeepSeek-V3 increases from 17.3% to 40.7%.
+These patterns persist on the after-knowledge-cutoff split; for example, DeepSeek-V3 improves from 44.3% to 67.9%, showing robustness of ECP when data contamination is controlled.
+On PutnamBench, absolute accuracies are lower due to the difficulty of the dataset, but ECP remains beneficial for all models.
+DeepSeek-V3.1 achieves state-of-the-art performance with 210 correct answer constructions via ECP, up from 181 problems with the CoT baseline.
+In addition, analysis of union accuracy shows that CoT and ECP correctly construct answers for overlapping but distinct subsets.
+For example, with DeepSeek-V3.1 as conjecturer, ECP can construct correct answers for 13.8% problems that CoT baseline fails.
+Still, 4.2% problems fail by ECP but successful with CoT baseline because ECP’s enumeration may under-generalize or fail on several problems, misleading the conjecturer in answer construction.
+Conjecturer Answer Construction End-to-End Problem Solving
+CoT / ECP Union CoT / ECP Union
+GPT-5 mini 69.7% / 73.6% 78.9% 32.5% / 33.1% 35.0%
+DeepSeek-V3 17.3% / 40.7% 43.4% 10.1% / 18.0% 18.7%
+DeepSeek-V3.1 65.4% / 75.0% 79.2% 31.1% / 32.7% 34.3%
+Table 2: Answer-construction accuracy and end-to-end problem solving accuracy of CoT baseline and ECP on ConstructiveBench (3640 problems in total).
+The first column shows the conjecturer models for answer construction, and the second column shows the answer-construction accuracies under CoT and ECP methods, along with their union accuracy.
+the third column shows end-to-end accuracies using DeepSeek-Prover-V2-7B for proof generation.
+Conjecturer Answer-Construction End-to-End Problem Solving
+CoT / ECP Union CoT / ECP Union
+GPT-5 mini 185 / 194 242 4 / 6 6
+DeepSeek-V3 64 / 107 124 1 / 3 3
+DeepSeek-V3.1 181 / 210 248 4 / 4 5
+Table 3: Number of correct answer constructions and end-to-end solved problems on PutnamBench (337 problems in total).
+The first column shows the conjecturer models for answer construction, with CoT and ECP results as well as their union.
+The second column shows the number of problems solved end-to-end with the DeepSeek-Prover-V2-7B prover (Pass@32).
+5.3 End-to-End Problem-Solving Accuracy
+In this section, we discuss the results of end-to-end problem solving across various combinations of conjecturer and prover models.
+The rightmost columns of Table 2 and Table 3 report accuracies using different conjecturer models for answer construction, combined with DeepSeek-Prover-V2-7B for proving.
+The complete full per-prover tables (including Goedel-Prover-SFT and Kimina-Prover-Preview-Distill-7B) for each dataset are placed in Appendix D.
+Overall, DeepSeek-Prover-V2-7B outperforms the other two prover models in these benchmarks.
+The strongest end-to-end problem-solving result on ConstructiveBench pairs GPT-5 mini as the conjecturer with DeepSeek-Prover-V2-7B as the prover, reaching 33.1% (up from 32.5% with CoT), with DeepSeek-V3.1 a close second at 32.7% (up from 31.1%).
+Even on the difficult PutnamBench answer-construction problems, ECP still solves 6 out of 337 problems end-to-end with GPT-5 mini and DeepSeek-Prover-V2-7B, compared to 4 out of 337 for CoT.
+The lower overall accuracy and smaller improvements are mainly due to two factors.
+First, incorrectly constructed answers lead to logically wrong statements that waste prover effort.
+Second, the prover models have limited power and often lack the capability to formally prove difficult theorems, especially ones in PutnamBench.
+Overall, ECP still provides consistent end-to-end gains over CoT, with the best-performing configuration given by GPT-5 mini as conjecturer and DeepSeek-Prover-V2-7B as prover.
 6 Limitations and Future Work
 Quality of Autoformalized Theorem Statements.
-Our framework and dataset depend on the quality of autoformalization produced by black-box LLMs.
+The ConstructiveBench depends on the quality of black-box LLM autoformalizations and LLM judge.
 In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics.
-While manual evaluation offers high fidelity, it is time-consuming and difficult to scale.
+While manual evaluation offers high fidelity and we evaluated the autoformalization quality of a subset of 100 problems, it is still time-consuming and intractable to large scale datasets.
 Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem.
@@
 Theorem-Proving.
-Our proving stage uses off-the-shelf theorem provers, which leaves significant room for optimization.
-In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime.
+State-of-the-art LLM is capable of winning gold medal in IMO exam (Huang and Yang, 2025), yet this achievement is out of reach for symbolic and LLM-based theorem provers in formal environment such as Lean.
+Our proving stage leverages off-the-shelf theorem provers, which still leaves significant room for optimization.
+Designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime.
 Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases.
 Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space.
-Fine-tuning LLMs or developing search-based algorithms for the formal answer-construction task thus represents a promising avenue for next-generation mathematical AI models.
 7 Conclusion
 We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean.
-To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedback.
-In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%.
-Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end.
-These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer-construction mathematics problems.
-Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.
+To support this work, we curated ConstructiveBench, a dataset of 3,640 answer‐construction problems formally verified through an iterative autoformalization pipeline with error feedback.
+In empirical evaluations, ECP improved the state of the art on ConstructiveBench, raising GPT-5 mini’s answer-construction accuracy from 69.7% to 73.6% and its end-to-end problem-solving accuracy from 32.5% to 33.1%.
+On the PutnamBench subset, ECP solves 6 out of 337 problems compared to 4 for the CoT baseline.
+These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in competition-level mathematical reasoning.
+Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to Olympiad and research-level mathematical problems.
 Appendix
@@
 Appendix B Prompt
-B.1 Prompt for Formalization
+B.1 Prompt for Autoformalization
+LLM Prompt You are given a math problem that requires answer-construction.
+Your task is to formalize the problem in Lean 4 (v4.23.0).
+1.
+Start with imports and namespaces such as `import Mathlib` and `open Real`.
+2.
+Define the ground-truth answer as an abbrev: `abbrev <problem_name>_answer : $\beta$ := <expression>` or `abbrev <problem_name>_answer : $\alpha$ $\to$ $\beta$ := fun x => <expression>` (Don't put `(n : $\mathbb{N}$)` before `: $\alpha$ $\to$ $\beta$`; encode inputs inside `fun x => ...`.) 3.
+Theorem skeleton (leave proof as sorry): `theorem <problem_name> (x : $\alpha$) (hypotheses about x) (y : $\beta$) : P(x, y) $\leftrightarrow$ y = <problem_name>_answer x := by sorry` Other Suggestions: - Use `IsLeast` / `IsGreatest` for optimal answers.
+- Use `Set $\beta$` and membership if there are multiple answers.
+B.2 Prompt for Enumerator
+LLM Prompt You will be given a math problem that requires answer construction in Lean.
+Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer.
+For problem with unfixed context parameters or infinite answers, try several parameters and bound the number of enumerations to 100.
+B.3 Prompt for Conjecturing
+LLM Prompt You will be given a math problem that requires answer construction in Lean.
+Your task is to reason about the problem and clearly state the closed-form final answer in Lean exprssion with explanations.
+You do not have to derive the whole proof.
+The program enumeration can be used as a hint.
+You must not "cheat" by an answer echoing problem statement.
+Appendix C Human Evaluation Examples
+To verify dataset quality and estimate the false-positive error rate, we randomly sampled 100 formalizations (from the 3,640 problems which have passed both the compiler and LLM judge) for manual evaluation, labeling each as: - Correct formalization - Formalization with minor error - Formalization with major error
+Below, we illustrate one example for each category. For clarity, we separate the answer and the statement, and use a placeholder for the answer variable in the Lean formalization.
+1. Correct formalization: omnimath19 (Chinese Mathematical Olympiad).
+Problem.
+Let $f:X\to X$, where $X=\{1,2,\ldots,100\}$, be a function satisfying: 1.
+$f(x)\neq x$ for all $x=1,2,\ldots,100$; 2.
+For any subset $A\subseteq X$ such that $|A|=40$, we have $A\cap f(A)\neq\emptyset$.
+Find the minimum $k$ such that for any such function $f$, there exists a subset $B\subseteq X$ with $|B|=k$ and $B\cup f(B)=X$.
+The answer is 69.
+Formalization.
+Listing 1: omnimath19 (correct formalization)
+import Mathlib
+open Set Function
+abbrev omnimath19_answer : $\mathbb{N}$ := 69
+theorem omnimath19 :
+IsLeast {k : $\mathbb{N}$ | 0 < k $\land$
+$\forall$ (f : Fin 100 $\to$ Fin 100),
+($\forall$ x, f x $\neq$ x) $\to$
+($\forall$ A : Finset (Fin 100), A.card = 40 $\to$ (A $\cap$ A.image f).Nonempty) $\to$
+$\exists$ B : Finset (Fin 100), B.card = k $\land$ B $\cup$ B.image f = Finset.univ }
+omnimath19_answer := by
+sorry
+Analysis.
+The constraint $X=\{1,\ldots,100\}$ is equivalently captured by Fin 100.
+The condition 1, $f(x)\neq x$, is formalized in ($\forall$ x, f x $\neq$ x).
+The constraint 2, $|A|=40$ and $A\cap f(A)\neq\emptyset$ is captured in using Finset.image, Finset.card and Nonempty syntax.
+The conditions regarding set $B$ is correctly formalized in similar manner.
+Finally, the target minimality condition is encoded using IsLeast.
+Overall, the formalization is correct without error.
+2. Formalization with minor error: omnimath1410 (HMMT).
+Problem. For positive integers $a,b$, let $M(a,b)=\frac{\mathrm{lcm}(a,b)}{\mathrm{gcd}(a,b)}$, and for $n\geq 2$ define
+$x_{n}=M(1,M(2,M(3,\ldots,M(n-2,M(n-1,n))\ldots))).$
+Compute the number of positive integers $n$ such that $2\leq n\leq 2021$ and $5x_{n}^{2}+5x_{n+1}^{2}=26x_{n}x_{n+1}$. The answer is 20.
+Listing 2: omnimath1410 (formalization with minor error)
+import Mathlib
+open Nat
+abbrev omnimath1410_answer : $\mathbb{N}$ := 20
+def M (a b : $\mathbb{N}$) : $\mathbb{N}$ := Nat.lcm a b / Nat.gcd a b
+def x : $\mathbb{N}$ $\to$ $\mathbb{N}$
+| 0 => 0
+| 1 => 1
+| n + 2 => M (n + 1) (x (n + 1)) -- off-by-one here on purpose
+theorem omnimath1410 :
+Finset.card
+(Finset.filter
+(fun n => 5 * (x n)^2 + 5 * (x (n + 1))^2 = 26 * x n * x (n + 1))
+(Finset.Icc 2 2021))
+= omnimath1410_answer := by
+sorry
+Analysis.
+The definition of $M$ and the theorem statement are correctly formalized, but the recursive definition of $x$ is off by one.
+Specifically, in this formalization $x_{2}=M(1,x(1))=1$, but the intended definition gives $x_{2}=M(1,2)=2$.
+This can be fixed by changing M (n + 1) (x (n + 1)) to M (n + 2) (x (n + 1)).
+3. Formalization with major error: IMO2023SLN7 (IMO).
+Problem. Let $a,b,c,d$ be positive integers satisfying
+$\frac{ab}{a+b}+\frac{cd}{c+d}=\frac{(a+b)(c+d)}{a+b+c+d}.$
+Determine all possible values of $a+b+c+d$. The answer is all positive integers that are not square-free.
+Listing 3: P2013AIMEII6 (formalization with major error)
+import Mathlib
+open Nat
+noncomputable abbrev IMO2023SLN7_answer : Set $\mathbb{N}$ := { n | $\exists$ p : $\mathbb{N}$, 2 $\leq$ p $\land$ p * p | n }
 […]
-‘‘‘lean
-your code
-B.2 Prompt for Enumerator Coding
-LLM Prompt I have a difficult high-school competition-level math problem.
-Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer.
-Use the following format for the delimited output:
-‘‘‘python
-your code
-B.3 Prompt for Conjecturing
-LLM Prompt I have a difficult high-school competition-level math problem in Lean, which asks to find a answer set with a mathematical proof.
-Your task is to clearly state the final answer in Lean exprssion with explanations.You do not have to derive the whole proof.
-I have enumerated the answer(s) via program as a hint for you, and you may look at the pattern in it to help propose the final answer.
-Your final answer should be in closed form in Lean and generalized from enumerated answer that can eliminate the existential quantifier.
-You must not echo the problem statement in your answer.
-Use the following format for the delimited output:
-‘‘‘lean
-your code
+Analysis.
+The logical shape is completely wrong.
+The problem asks for all possible values of $a+b+c+d$ that occur for some positive integers satisfying the equation.
+However, the formalized theorem states that for some specific $a,b,c,d>0$ the equation holds iff $a+b+c+d\in answer$, which would mean the equation is true for those exact $a,b,c,d$ precisely when the sum is not square-free.
+Appendix D Full Tables
+D.1 ConstructiveBench Full (3640 problems)
+DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
+Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
+GPT-5 nano 28.8% 30.9% 34.5% 24.9% 26.1% 29.5% 22.3% 23.2% 27.4%
+GPT-5 mini 32.5% 33.1% 35.0% 27.2% 27.6% 29.1% 25.0% 25.2% 27.6%
+DeepSeek-V3 10.1% 18.0% 18.7% 7.0% 15.0% 15.3% 7.8% 14.0% 14.8%
+DeepSeek-V3.1 31.1% 32.7% 34.3% 26.2% 27.5% 29.0% 23.8% 25.1% 27.6%
+GPT-4.1 mini 7.3% 17.0% 17.7% 5.3% 11.1% 12.0% 6.0% 13.1% 13.9%
+GPT-4.1 nano 9.8% 14.3% 16.7% 7.1% 10.0% 11.6% 7.9% 11.2% 13.1%
+Table 4: ConstructiveBench full end-to-end results (Pass@32)
+D.2 ConstructiveBench After-Cutoff (106 problems)
+DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
+Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
+GPT-5 nano 13.2% 15.1% 18.9% 11.3% 12.3% 14.2% 10.4% 11.3% 13.2%
+GPT-5 mini 17.9% 22.6% 22.6% 15.1% 16.0% 17.9% 13.2% 14.2% 15.1%
+DeepSeek-V3 2.8% 5.7% 5.7% 1.9% 4.7% 4.7% 1.9% 3.8% 3.8%
+DeepSeek-V3.1 17.0% 17.9% 20.8% 15.1% 16.0% 17.9% 13.2% 14.2% 15.1%
+GPT-4.1 mini 4.7% 6.6% 6.6% 3.8% 5.7% 5.7% 3.8% 4.7% 4.7%
+GPT-4.1 nano 2.8% 4.7% 5.7% 1.9% 2.8% 3.8% 1.9% 2.8% 3.8%
+Table 5: ConstructiveBench After-Cutoff end-to-end results (Pass@32)
+D.3 PutnamBench Subset (337 problems)
+DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
+Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
+GPT-5 nano 1.2% 1.5% 1.8% 0.9% 0.9% 1.2% 0.9% 1.2% 1.5%
+GPT-5 mini 1.2% 1.8% 1.8% 0.9% 0.3% 0.9% 0.9% 0.6% 0.9%
+DeepSeek-V3 0.3% 0.9% 0.9% 0.0% 0.3% 0.3% 0.0% 0.0% 0.0%
+DeepSeek-V3.1 1.2% 1.2% 1.5% 0.6% 0.3% 0.6% 0.6% 0.6% 0.9%
+GPT-4.1 mini 0.6% 0.6% 0.9% 0.6% 0.6% 0.9% 0.0% 0.0% 0.0%
+GPT-4.1 nano 0.9% 0.6% 0.9% 0.3% 0.0% 0.3% 0.0% 0.0% 0.0%
+Table 6: PutnamBench subset end-to-end results (Pass@32) with timeout=120s.
+D.4 Full Answer-Construction Accuracies
+Conjecturer ConstructiveBench (Full) ConstructiveBench (After-Cutoff) PutnamBench
+CoT / ECP Union CoT / ECP Union CoT / ECP Union
+GPT-5 nano 45.1% / 62.3% 66.9% 28.3% / 55.7% 57.6% 16.3% / 32.3% 36.8%
+GPT-5 mini 69.7% / 73.6% 78.9% 51.9% / 58.5% 64.2% 54.9% / 57.6% 71.8%
+DeepSeek-V3 17.3% / 40.7% 43.4% 2.8% / 26.4% 26.4% 19.0% / 31.8% 36.8%
+DeepSeek-V3.1 65.4% / 75.0% 79.2% 44.3% / 67.9% 69.8% 53.7% / 62.3% 73.6%
+GPT-4.1 mini 16.6% / 36.5% 37.8% 11.3% / 23.6% 24.5% 15.1% / 23.2% 29.4%
+GPT-4.1 nano 15.3% / 28.5% 33.0% 4.7% / 18.9% 20.8% 7.7% / 11.6% 17.2%
+Table 7: Answer-construction accuracy. Note that, DeepSeek-V3.1 has knowledge cutoff after June 2024, so this model’s ConstructiveBench (After-Cutoff) accuracy may be contaminated.